OpenStudy (anonymous):

evaluate the limit as x goes to 0: 6x/tan(8x)?

OpenStudy (amistre64):

lhopital it; and see what you get

OpenStudy (amistre64):

6/8sec^2(8x)

OpenStudy (anonymous):

thanks, no t allowed to use lhopital :-)

OpenStudy (amistre64):

(6/8) cos^2(0) = 3/4

OpenStudy (anonymous):

yay! that's what I got!!!

OpenStudy (amistre64):

not allowed to use ?? lol. the answer doesnt really care how its found

OpenStudy (anonymous):

haha, I hear you> I asked my teacher abt L'hopital and she was very adamant taht I show calc's the way she is teaching us :-(

OpenStudy (amistre64):

i want you to move 3 feet to your left; but in order to get there, you have to run around the block first ...lol

OpenStudy (amistre64):

the other way i spose is just throw in the f(x+h)-f(x)/h conflageration

OpenStudy (anonymous):

I completely agree with you!

OpenStudy (anonymous):

we have to sub sin and cos and then use the rules

myininaya (myininaya):

\[\frac{6x}{\frac{\sin(8x)}{\cos(8x)}}=6x*\frac{\cos(8x)}{\sin(8x)}=6x*\cos(8x)*\frac{1}{\sin(8x)}\] =\[6*\cos(8x)*\frac{8x}{\sin(8x)}*\frac{1}{8}\]

OpenStudy (anonymous):

just put another one up

myininaya (myininaya):

x->0, cos(8x)->1 x->0, 8x/sin(8x)->1 so we have 6*1/8=6/8=3/4

OpenStudy (anonymous):

thanks Myin, I got it though